if the side of triangle are 16,20 and 27 respectivly,what is the lenght of the bisector of the largest angle?


1
Expert's answer
2021-12-20T16:38:04-0500






Using cosine rule;


"27^2=16^2+20^2-2×16×20\>cos \>2\theta"


"2\theta=96.55"

"\theta=48.28"


"20^2=27^2+16^2-2×16×27\>Cos\>\beta"


"\beta=47.38"


"\frac{16}{Sin\>84.34}=\frac{x}{Sin\>47.38}"


"x=11.83"